OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 3x - y = -3 and passes through the point (4, -2). help me pleaseeeee

OpenStudy (anonymous):

if you wont answer i cant help you :C

OpenStudy (anonymous):

1) y = - 1/3 x - 6 2) y = -1/3 x - 14 3) y = 3x - 6 4) y = 3x - 14

OpenStudy (jdoe0001):

can you solve " 3x - y = -3 " for "y"? what would that give you?

OpenStudy (anonymous):

im sorry I just made one of these and im not sure how to work it

OpenStudy (jdoe0001):

http://www.youtube.com/watch?v=n3FbQGGKVfc <- you may want to brush up your linear equation material then

OpenStudy (anonymous):

am I solving for y

OpenStudy (precal):

yes solve for y, then look at the number in front of x, that is your slope value. Parallel lines have the same slope value.